demand and supply equation for a perfect competition market are given as Qd=100-3P and Q's= 40 +P what will the firm's in that market take

1
Expert's answer
2021-05-18T09:40:06-0400

At a perfectly competitive market, "Qd=Qs" at equilibrium.

"Qd=100-3P"

"Qs=40+P"


Therefore, "100-3P=40+P"

"100-40=P+3P"

"60=4P"

"P=15"


"Qs=40+P"

"40+15=55"

"Q=55" units.